OpenStudy (anonymous):

a force of 5n is applied on a body of mass 20 kg. Calculate the work done in 3rd second.

OpenStudy (anonymous):

F=ma a = F/m Distance Travelled in nth Second is Dn = u + a/2 (2n-1) Assumiing the Body starts From rest u = 0 W = F * Dn hope that helps..)